McHugh Company allocates common Building Department costs to producing departments (P1 and P2) based on space occupied, and it allocates common Personnel Department costs based on the number of employees. Space occupancy and employee data are as follows:

Building Personnel Dept. P1 Dept. P2
Space occupied 2,000 ft. 10,000 ft. 120,000 ft. 70,000 ft.
Employees 6 10 80 50


If McHugh Company uses the sequential allocation method and the support department with the highest percentage of interdepartmental services is allocated first, the ratio representing the portion of Personnel Department costs allocated to Department P2 is

a.90/140.

b.50/140.

c.50/130.

d.50/146.
Answer #1

Number of employees in total = 80 + 50 = 130

P2 department employees = 50

Personnel dept costs allocated = P2 dept employees / Total

= 50/130

Option C is the answer